# Enhancing Autonomous Navigation by Imaging Hidden Objects using Single-Photon LiDAR

### v1

@@ I Introduction
Robust autonomous navigation remains a critical challenge in robotics as mobile robots become increasingly prevalent, from industrial warehouses to urban environments. Robust navigation systems have the potential to save lives by preventing accidents, reduce financial losses by protecting human and robotic assets, and enhance operational efficiency by optimizing operation speed and energy consumption.
@@
We evaluated the effectiveness of our NLOS perception approach for autonomous navigation through simulation and real-world experiments. We develop a simulation framework that couples transient rendering with vehicle dynamics within the multi-physics engine Project Chrono[1][2]. This allows us to co-simulate and model the multi-bounce histograms of single-photon LiDARs with vehicle/robot dynamics simulations. Our simulation experiments demonstrate NLOS perception aids in navigating an L-turn corner with reduced chances of collision with hidden obstacles compared to LOS-only perception. We implemented our system on a physical two-wheeled robot equipped with a commercial-grade single-photon LiDAR. In a blind corner navigation task, the NLOS-capable robot demonstrated superior performance, generating and following more efficient paths around hidden objects. Quantitatively, the NLOS approach reduced travel time by more than half and shortened the trajectory by 33% compared to LOS-only navigation.
@@ II-A Single-photon LiDAR
Single-photon LiDAR has emerged as a promising sensing technology for robotics applications. Unlike traditional LiDAR systems that typically record only the first return, single-photon LiDARs can detect and time-stamp individual photons with picosecond-level precision [3, 4]. Single-photon LiDAR has higher sensitivity and allows one to capture not just the direct reflections from visible surfaces, but also the indirect reflections that have undergone multiple bounces before returning to the sensor. Single-photon LiDAR has shown great promise in depth sensing [5, 6], 3D surface mapping [7, 8], handling specular surfaces [9, 10], long-range detection [11, 12], airborne applications such as snow depth mapping [13, 14, 9], forest surveying [15], and 3D terrain mapping [16].
@@ II-B Non-Line-Of-Sight Imaging
Non-line-of-sight imaging has emerged as a promising application by utilizing single-photon LiDAR that can measure indirect reflections and image objects hidden from the direct line of sight. Recent works demonstrate 3D reconstruction of occluded scenes from single-photon LiDAR at high spatial resolution and computational efficiency [19, 20, 21, 22, 23, 24]. While the majority of existing works focus on 3D imaging, there are also notable contributions to 3D pose estimation [25, 26] and 3D tracking [4] using NLOS information. Our paper presents the first experimental demonstration of NLOS imaging for robotics applications, paving the way for novel applications in autonomous navigation and sensing.
@@ III-A Sensing histograms using single-photon LiDAR
We utilize Flash single-photon LiDARs [17, 18], which comprises of an array of SPAD pixels that each generates its own time-of-flight histogram. In this paper, we consider that the SPAD LiDAR is pointed at a planar floor, and that there is single hidden object of unknown location $(x_{h},y_{h})$ and size $s_{h}$. The spatio-temporal histogram array obtained $\mathbf{H}$ depends on $x_{h}$, $y_{h}$ and $s_{h}$.
@@ III-B Perception of the hidden region
Estimating size and location of a hidden object from the histogram array is challenging. The first challenge is tied to the ambiguity of the signal signature. There could be multiple hidden object positions that could have the same 3-bounce time-of-flight. The second challenge is that of low signal quality. Commercial LiDARs operate with eye-safe lasers imposing a limit on the emitted laser power. Three bounces on surfaces severely attenuates the signal reaching back to the sensor resulting in very low signal-to-noise ratio (SNR).
We tackle the ambiguity and SNR challenges of commercial-grade LiDAR sensing using a data-driven approach. We train a convolutional neural network (CNN) that inputs the histogram array and outputs the size and location of the hidden object. This network is trained on a dataset of histogram arrays captured with varying hidden object location and size along with the LiDAR location.
$\left(\hat{x}_{h},\hat{y}_{h},\hat{s}_{h}\right)=f_{\theta}\left(\mathbf{H}\left(x_{h},y_{h},s_{h}\right)\right)$ (1)
@@ IV Simulation Results
Figure 3: Simulation Results for navigating a L-turn corner using line-of-sight and non-line-of-sight perception. a) Comparison of collision avoidance success rate between LOS and NLOS perception. NLOS has an edge in performance over LOS at higher speeds where path planning cannot react on time. b) And, we find that higher localization error leads to worse performance in avoiding obstacles. Which matches intuition.
@@ IV-A Dynamics-Integrated Transient Rendering
To design robotic controllers that utilize transient data in-the-loop, we have developed a transient renderer within the open-source Chrono multi-physics engine [1]. With native integration of other sensors [2], vehicles [38], and ROS [39, 40], Chrono enables the simulation of a wide range of robotic applications. Transient rendering is the simulation of the propagation of light at very high time resolutions. As opposed to standard path tracing, transient rendering models the light at a point in the spatial and time domains ($x,y,t$). Using transient rendering, we can model the amount of light arriving at a time-of-flight detector often at time resolutions on the order of picoseconds or less and record it as a collection of histograms for each time interval. Using these high-time-resolution histograms, we can distinguish light coming from multiple scattering/bounce events in the scene. This enables us to infer information about the scene even beyond the line of sight. Our transient renderer is based on Royo et al.’s work on non-line-of-sight transient rendering [41]. Since NLOS transient rendering is computationally expensive, we use GPU hardware accelerated ray tracing and special sampling methods [41] to improve rendering speeds. Our simulation framework can model vehicle dynamics and render an $8\times 8$ multi-bounce transient sensor at 10 FPS.
IV-B High-speed Hallway Navigation with Occluded Obstacles
We view the use of multi-bounce light as an enhancement to existing perception-based approaches for robotic navigation. Our simulated experiments, shown in Figure 1, deploy a wheeled vehicle in a corridor environment with a hidden retro-reflective object of diameter of 2m randomly placed around the corner. The vehicle maintains a constant speed throughout the maneuver. The vehicle is then either provided sensor data from a single-bounce LiDAR (Figure 1a), or from a multi-bounce LiDAR (Figure 1b). The perceived area is conceptualized as an occupancy map, where obstacles and walls are considered “occupied” and everywhere else is “unoccupied”. Training data for the NLOS object detection model is acquired by running the vehicle using a privileged controller through 30 different environments with unique object positions and saving the transient information. With the trained model, we run a total of 200 trials with uniformly sampled speed values for both LOS and NLOS perception. Analysis of the NLOS vs LOS-based navigation is shown in Figure 3a.
@@ IV-B High-speed Hallway Navigation with Occluded Obstacles
where $r_{max}$ is the maximum turn radius of the corner ($15m$), $mu$ is the coefficient of friction of the vehicle’s tires (estimated to be $0.9$), and $g$ is acceleration due to gravity. As such, for our simulation experiments we set $12ms^{-1}$ as the maximum velocity of the vehicle.
@@ V-A Experimental Setup
Our experimental setup is composed of a two wheeled mobile robot navigation around an L-shaped corridor. We utilize the AMS TMF8828 SPAD-based LiDAR module and their Arduino evaluation kit as the sensor system [17]. The TMF8828 sensor is comprised of $18\times 12$ individual SPAD pixels, which is aggregated into $3\times 3$ measurement zones. In this configuration, the sensor has an effective FoV of $41^{\circ}\times 52^{\circ}$ and can measure light traveling a total of 10 m. Ground truth position measurements are obtained via four ArUco tags [42] – placed on the scene, robot, object, and goal respectively – and imaged with a monochrome camera mounted with a birds-eye view of the entire scene. The robot is a small, two wheeled mobile robot [43] equipped with a Raspberry Pi 5 for processing. Training of the underlying perception model is performed on an external system, but data acquisition and evaluation takes place on the robot with ground truth positions streamed over the network.
@@ V-B Data Acquisition and Training
Figure 4: Capture and evaluation setup. a) To validate the perception algorithm and generate a dataset to train the CNN, we utilize a 2-axis gantry to precisely position a hidden object (in this case, a retro-reflective arrow) in some region outside the FoV of the SPAD sensor. For training, we generate data for many positions and orientations of both the sensor and object to produce a diverse dataset. b) Multi-bounce histograms of four of the output zones from the SPAD sensor. The first bounce is the primary spike, and in later bins (i.e. with a longer time-of-flight), the third bounce reflection from the object can be observed. Different zones have a different view of the scene, and so the relative intensity and peak position encode the object’s state. This information is leveraged by the model to predict the position of the hidden object. c) The estimates of our model on an evaluation set (i.e., data held out from training) which shows our model estimates object position with sub-10 cm accuracy. We found that evaluation performance decreased substantially when testing on states (i.e., camera/object positions) not seen during training.
@@ V-C Blind Corner Navigation
Figure 5: Blind corner navigation using non-line-of-sight perception on a physical robot. To test the effectiveness of non-line-of-sight imaging for autonomous navigation, we ran experiments on a physical robot in an L-shaped blind corner. An object was placed in an area hidden from view of the sensor. a) The agent with only a LOS view couldn’t prepare for the object before rounding the corner, and had to overshoot back inside to reach the goal. b) On the other hand, when NLOS information was used, the robot could successfully generate and follow a path which reduced the wasted travel time. c) A quantitative depiction of the efficiency improvements for an NLOS-capable robot. With only LOS information, the robot took more than 2$\times$ longer and required a 33% longer trajectory.

### v2

@@ I Introduction
Robust autonomous navigation remains a critical challenge in robotics as mobile robots become increasingly prevalent, from industrial warehouses to urban environments. These systems have the potential to save lives by preventing accidents, reduce financial losses by protecting human and robotic assets, and enhance operational efficiency by optimizing operation speed and energy consumption.
@@
We evaluated the effectiveness of our NLOS perception approach for autonomous navigation through simulation and real-world experiments. We develop a simulation framework that couples transient rendering with vehicle dynamics within the multi-physics engine Project Chrono [1, 2]. This allows us to co-simulate and model the multi-bounce histograms of single-photon LiDARs with vehicle/robot dynamics simulations. Our simulation experiments demonstrate NLOS perception aids in navigating an L-turn corner with reduced chances of collision with hidden obstacles compared to LOS-only perception. We implemented our system on a physical two-wheeled robot equipped with a commercial-grade single-photon LiDAR. In a blind corner navigation task, the NLOS-capable robot demonstrated superior performance, generating and following more efficient paths around hidden objects. Quantitatively, the NLOS approach reduced travel time by more than half and shortened the trajectory by 33% compared to LOS-only navigation.
@@ II-A Single-photon LiDAR
Single-photon LiDAR has emerged as a promising sensing technology for robotics applications. Unlike traditional LiDAR systems that typically record only the first return, single-photon LiDARs can detect and time-stamp individual photons with picosecond-level precision [3, 4]. Single-photon LiDAR has higher sensitivity and allows one to capture not just the direct reflections from visible surfaces, but also the indirect reflections that have undergone multiple bounces before returning to the sensor. Single-photon LiDAR has shown great promise in depth sensing [5, 6], 3D surface mapping [7, 8, 9], handling specular surfaces [10, 11], long-range detection [12, 13], airborne applications such as snow depth mapping [14, 15, 10], forest surveying [16], and 3D terrain mapping [17].
@@ II-B Non-Line-Of-Sight Imaging
Non-line-of-sight imaging has emerged as a promising application by utilizing single-photon LiDAR that can measure indirect reflections and image objects hidden from the direct line of sight. Recent works demonstrate 3D reconstruction of occluded scenes from single-photon LiDAR at high spatial resolution and computational efficiency [20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30]. While the majority of existing works focus on 3D imaging, there are also notable contributions to 3D pose estimation [31, 32] and 3D tracking [4] using NLOS information. However, these works typically rely on expensive research-grade sensors. Only a few works have shown NLOS imaging with commercial-grade sensors [4]. Our paper presents the first experimental demonstration of NLOS imaging for robotics applications, paving the way for novel applications in autonomous navigation and sensing.
@@ III-A Sensing histograms using single-photon LiDAR
We utilize Flash single-photon LiDARs [18, 19], which comprises of an array of SPAD pixels that each generates its own time-of-flight histogram. In this paper, we consider that the SPAD LiDAR is pointed at a planar floor, and that there is single hidden object of unknown location $(x_{h},y_{h})$. The spatio-temporal histogram array obtained $\mathbf{H}$ thus depends on $x_{h}$ and $y_{h}$.
$\left(\hat{x}_{h},\hat{y}_{h}\right)=f_{\theta}\left(\mathbf{H}\left(x_{h},y_{h}\right)\right)$ (1)
@@ III-B Perception of the hidden region
Estimating location of a hidden object from the histogram array is challenging. The first challenge is tied to the ambiguity of the signal signature. There could be multiple hidden object positions that have the same 3-bounce time-of-flight. The second challenge is that of low signal quality. Commercial LiDARs operate with eye-safe lasers imposing a limit on the emitted laser power. Three bounces on surfaces severely attenuates the signal reaching back to the sensor resulting in very low signal-to-noise ratio (SNR).
We tackle the ambiguity and SNR challenges of commercial-grade LiDAR sensing using a data-driven approach. We train a convolutional neural network (CNN) that inputs the histogram array and outputs the location of the hidden object. This network is trained on a dataset of histogram arrays captured by a LiDAR with at different positions with varying hidden object locations and sizes.
@@ IV Simulation Results
Figure 3: Simulation Results for navigating a L-turn corner using line-of-sight and non-line-of-sight perception. (a) Comparison of collision avoidance success rate between LOS and NLOS perception. NLOS has an edge in performance over LOS at higher speeds where path planning cannot react on time. (b) And, we find that higher localization error leads to worse performance in avoiding obstacles. Which matches intuition.
@@ IV-A Dynamics-Integrated Transient Rendering
To design robotic controllers that utilize transient data in-the-loop, we have developed a transient renderer within the open-source Chrono multi-physics engine [1]. With native integration of other sensors [2], vehicles [44], and ROS [45, 46], Chrono enables the simulation of a wide range of robotic applications. Transient rendering is the simulation of the propagation of light at very high time resolutions. As opposed to standard path tracing, transient rendering models the light at a point in the spatial and time domains ($x,y,t$). Using transient rendering, we can model the amount of light arriving at a time-of-flight detector often at time resolutions on the order of picoseconds or less and record it as a collection of histograms for each time interval. Using these high-time-resolution histograms, we can distinguish light coming from multiple scattering/bounce events in the scene. This enables us to infer information about the scene even beyond the line of sight. Our transient renderer is based on the non-line of sight transient rendering introduced by Royo et al. [47]. Since NLOS transient rendering is computationally expensive, we use GPU hardware accelerated ray tracing and special sampling methods [47] to improve rendering speeds. Our simulation framework can model vehicle dynamics and render an $8\times 8$ multi-bounce transient sensor at 10 FPS.
IV-B High-speed Blind Corner Navigation with Occluded Obstacles
We view the use of multi-bounce light as an enhancement to existing perception-based approaches for robotic navigation. Our simulated experiments, shown in Figure 1, deploy a full-scale wheeled vehicle in a cityscape environment with a randomly placed hidden cyclist wearing retro-reflective gear. We approximate the cyclist with a 3D bounding volume of length 2m. The width and height of the simulation scene is approximately 50m. The vehicle maintains a constant speed throughout the maneuver. The vehicle is then either provided sensor data from a single-bounce LiDAR (Figure 1a), or from a multi-bounce LiDAR (Figure 1b). The perceived area is conceptualized as an occupancy map, where obstacles and walls are considered “occupied” and everywhere else is “unoccupied”. Training data for the NLOS object detection model is acquired by running the vehicle using a controller with privileged information on obstacle positions through 30 different environments with unique object positions and saving the transient information. With the trained model, we run a total of 200 trials with uniformly sampled speed values for both LOS and NLOS perception. Analysis of this NLOS vs LOS-based navigation is shown in Figure 3a.
@@ IV-B High-speed Blind Corner Navigation with Occluded Obstacles
where $r_{max}$ is the maximum turn radius of the corner ($15m$), $\mu$ is the coefficient of friction of the vehicle’s tires (estimated to be $0.9$), and $g$ is acceleration due to gravity. As such, for our simulation experiments we set $12ms^{-1}$ as the maximum velocity of the vehicle.
@@ V-A Experimental Setup
Our experimental setup is composed of a two wheeled mobile robot navigation around an L-shaped corridor. We utilize the AMS TMF8828 SPAD-based LiDAR module and their Arduino evaluation kit as the sensor system [18]. The TMF8828 sensor is comprised of $18\times 12$ individual SPAD pixels, which is aggregated into $3\times 3$ measurement zones. In this configuration, the sensor has an effective FoV of $41^{\circ}\times 52^{\circ}$ and can measure light traveling a total of 10 m. We acquire sensor information at roughly 5 Hz. Ground truth position measurements are obtained via four ArUco tags [48] – placed on the scene, robot, object, and goal respectively – and imaged with a monochrome camera mounted with a birds-eye view of the entire scene. The robot is a small, two wheeled mobile robot [49] equipped with a Raspberry Pi 5 for processing. Training of the underlying perception model is performed on an external system, but data acquisition and evaluation takes place on the robot with ground truth positions streamed over a network. Inference of the network on the Raspberry Pi 5 takes approximately 50 milliseconds.
@@ V-B Data Acquisition and Training
Figure 4: Capture and evaluation setup. (a) To validate the perception algorithm and generate a dataset to train the CNN, we utilize a 2-axis gantry to precisely position a hidden object (in this case, a retro-reflective arrow) in some region outside the FoV of the SPAD sensor. For training, we generate data for many positions and orientations of both the sensor and object to produce a diverse dataset. (b) Multi-bounce histograms of four of the output zones from the SPAD sensor. The first bounce is the primary spike, and in later bins (i.e. with a longer time-of-flight), the third bounce reflection from the object can be observed. Different zones have a different view of the scene, and so the relative intensity and peak position encode the object’s state. This information is leveraged by the model to predict the position of the hidden object. (c) The estimates of our model on an evaluation set (i.e., data held out from training) which shows our model estimates object position with sub-10 cm accuracy. We found that evaluation performance decreased substantially when testing on states (i.e., camera/object positions) not seen during training.
@@ V-C Blind Corner Navigation
Figure 5: Blind corner navigation using non-line-of-sight perception on a physical robot. To test the effectiveness of non-line-of-sight imaging for autonomous navigation, we ran experiments on a physical robot in an L-shaped blind corner. An object was placed in an area hidden from view of the sensor. (a) The agent with only a LOS view couldn’t prepare for the object before rounding the corner, and had to overshoot back inside to reach the goal. (b) On the other hand, when NLOS information was used, the robot could successfully generate and follow a path which reduced the wasted travel time. (c) A quantitative depiction of the efficiency improvements for an NLOS-capable robot. With only LOS information, the robot took more than 2$\times$ longer and required a 33% longer trajectory.